OpenStudy (anonymous):

Indicate the equation of the given line in standard form. The line with slope 9/7and containing the midpoint of the segment whose endpoints are (2, -3) and (-6, 5).

jimthompson5910 (jim_thompson5910):

What is the midpoint of (2, -3) and (-6, 5) ?

OpenStudy (anonymous):

9/7 ?

jimthompson5910 (jim_thompson5910):

That's the given slope

jimthompson5910 (jim_thompson5910):

You need to use the midpoint formula. Are you familiar with that formula?

OpenStudy (anonymous):

then i dont know ?

jimthompson5910 (jim_thompson5910):

The midpoint formula is http://geometry.jdeer.com/Images/c1/midpoint%20formula.gif

jimthompson5910 (jim_thompson5910):

Essentially what you do is add up the corresponding coordinates. Then divide by 2.

OpenStudy (anonymous):

I know the formela but i get confused on how to use it its really hard

jimthompson5910 (jim_thompson5910):

For instance, add up the x coordinates to get 2 + (-6) = 2 - 6 = -4 Then divide that by 2 to get -4/2 = -2 So the x coordinate of the midpoint is x = -2

jimthompson5910 (jim_thompson5910):

What do you get when you do that with the y coordinates?

OpenStudy (anonymous):

m=-3-5=4-5

OpenStudy (anonymous):

like that because i dont know its to confusing

jimthompson5910 (jim_thompson5910):

actually it should be -3 + 5 = 2 Then you divide by 2 to get 2/2 = 1 So the y coordinate of the midpoint is y = 1

jimthompson5910 (jim_thompson5910):

Therefore, the midpoint is (-2,1)

jimthompson5910 (jim_thompson5910):

So what you need to do now is find the equation of the line with slope = 9/7 and it goes through the point (-2,1)

OpenStudy (anonymous):

do I divide 9/7 and -2 1

jimthompson5910 (jim_thompson5910):

no you must use y = mx+b and plug in m = 9/7, x = -2, y = 1 Then solve for b

OpenStudy (anonymous):

i put it into a caulator and i got -126

jimthompson5910 (jim_thompson5910):

y = mx+b 1 = (9/7)*(-2)+b 1 = -18/7+b solve for b

OpenStudy (anonymous):

i dont know how to make the eqaution for the answer

jimthompson5910 (jim_thompson5910):

You need to add 18/7 to both sides

OpenStudy (anonymous):

i dont know if this is correct but 41 and 27

jimthompson5910 (jim_thompson5910):

nope

OpenStudy (anonymous):

46

OpenStudy (anonymous):

so what does that eqaution look like its mx+b=c right

jimthompson5910 (jim_thompson5910):

1 + 18/7 7/7 + 18/7 (7+18)/7 25/7

jimthompson5910 (jim_thompson5910):

So, 1 + 18/7 = 25/7 which means b = 25/7

jimthompson5910 (jim_thompson5910):

m = 9/7 (given) b = 25/7 (just found this)

jimthompson5910 (jim_thompson5910):

plug those into y = mx+b to get the equation

OpenStudy (anonymous):

ok so x+9/7=25/7

OpenStudy (anonymous):

like that ?....

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

See attached

OpenStudy (anonymous):

x Y - - - - - - - - - -

OpenStudy (anonymous):

okay i understand that but how to i plug it in to my answer so it makes sence

jimthompson5910 (jim_thompson5910):

Are you able to post a screenshot of that?

OpenStudy (anonymous):

sorry its not letting me

OpenStudy (anonymous):

But it is what it looks like i just have to put the numbers in so they fit with x and y

jimthompson5910 (jim_thompson5910):

Oh you just have to make the equation look like what I posted in the final answer

jimthompson5910 (jim_thompson5910):

So you have to create a fraction 9/7 for the slope and 25/7 for the y-intercept

OpenStudy (anonymous):

SWEET THANKS!:) for are your help :)

jimthompson5910 (jim_thompson5910):

you're welcome